OpenStudy (anonymous):

log x - log (x-3) = 1 Do not state any extraneous roots. If the answer is not exact, round to four decimal places.

OpenStudy (dumbcow):

log a - log b = log (a/b) i am assuming this is log base 10 ? \[\log (\frac{x}{x-3}) = 1\] \[\frac{x}{x-3} = 10\] solve for x